# thanks: corresponding author

### v1

@@ 4.1 Exceptions for SVM Kernels
Handling Non-PSD Kernels in SVMs
To address the limitations of non-PSD kernels, several strategies have been proposed, which include
  • Spectral correction methods: involve clipping, flipping, or shifting eigenvalues. These techniques can modify the kernel matrix to enforce positive semi-definiteness. Despite their simplicity, spectral correction methods can result in the loss of crucial structural information embedded in the original kernel [16, 23].
  • Pseudo-Euclidean embedding: Embedding the data into Krein spaces or pseudo-Euclidean spaces has been explored as a way to handle indefinite kernels. Although effective, these methods increase computational complexity and are less interpretable [21, 15].
  • Difference of Convex Functions (DC) programming: SVM optimization problems can be reformulated using DC programming into convex and concave parts. This approach ensures convergence to a locally optimal solution but may involve higher computational overhead [17].
  • Iterative methods: iterative techniques, including Sequential Minimal Optimization (SMO) [25], have been adapted for training SVMs with non-PSD kernels. These methods introduce stability improvements while maintaining efficiency [24].
As indicated above, each of these strategies has its own limitations, which warrants the development of other alternative approaches. Here, we propose a custom kernel construction based on the Kernel Polynomial Method. Such a custom kernel can be used whenever there is no close correspondence between a kernel and a Green’s function.
KPM-Based Kernel Construction
The KPM framework approximates a given kernel $K(x,y)$ through a series expansion:
$K(x,y)\approx\sum_{k=0}^{n}c_{k}T_{k}(x^{\top}y),$ (30)
where $T_{k}$ are orthogonal polynomials, and the coefficients $c_{k}\geq 0$ are chosen to enforce positivity. This expansion aligns with Mercer’s theorem, ensuring that the resulting kernel satisfies the necessary positive semi-definiteness conditions required for SVM optimization [13, 2].
Unlike spectral correction methods, which may lose critical structural information in the kernel matrix, KPM-based kernels retain the original kernel’s functional form while ensuring PSD properties. By selecting specific polynomial bases (e.g., Chebyshev or Jacobi polynomials), KPM can adapt to various kernel properties and target applications [27]. Furthermore, KPM naturally aligns with the mathematical structure of Green’s functions, which frequently appear in physical systems to model electron propagation and Gaussian decay [19]. This connection makes KPM-based kernels particularly suitable for applications involving physical models, such as quantum mechanics or biological similarity measurements [24].
4.2 Incorporation of the Kernel Polynomial Method (KPM)
Traditionally, KPM has relied on Chebyshev polynomials for their optimal convergence properties on finite intervals. Recent developments, however, show that Jacobi polynomials [28] can extend the flexibility of KPM by incorporating additional parameters that adjust the weight functions, enabling tailored kernel designs for specific spectral properties.[28] These expansions provide both theoretical guarantees, such as positivity and Mercer compliance, and practical algorithms for constructing kernels that align with quantum Green’s functions. This is particularly advantageous when handling systems where Gaussian-like decay and spectral accuracy are critical. One could, in principle, employ other families of orthogonal polynomials, such as Jacobi polynomials, to incorporate custom weight functions, although explicit demonstrations of such expansions remain limited in the literature. For instance, it has been demonstrated that by leveraging orthogonal polynomial expansions, KPM enables efficient approximations of Green’s functions associated with physical operators, such as Hamiltonians in quantum systems [26, 27, 28].
4.2.1 Core Concepts of KPM
@@
where $T_{n}(x)$ are Chebyshev polynomials, and $\mu_{n}$ are expansion coefficients. To ensure stability and convergence, the spectrum of $A$ is scaled to $[-1,1]$ and damping kernels (e.g., Jackson kernel) adjust the coefficients:
@@
These steps guarantee that the resulting kernel remains positive semidefinite and closely approximates the desired Green’s function.
This method parallels the representer theorem, as both yield solutions expressed as finite linear combinations of kernel (or Green’s function) evaluations. Mercer’s theorem ensures that kernels constructed through such polynomial approximations remain positive semidefinite, facilitating valid SVM optimization. By approximating Green’s functions with Chebyshev expansions, KPM also bridges theoretical operator inversion with practical kernel design, grounding machine learning models in physical reality while maintaining rigorous mathematical properties.
Using KPM, one can design custom kernels for specific physical problems. For instance, to model the density of states (DOS) in a quantum system, KPM approximates the spectral function related to the Green’s function:
$K(\mathbf{x},\mathbf{x}^{\prime})=\sum_{n=0}^{N}\mu_{n}\,T_{n}\bigl{(}\gamma\,\mathbf{x}^{\top}\mathbf{x}^{\prime}+r\bigr{)},$
where coefficients $\mu_{n}$ incorporate kernel weights that ensure positive semidefiniteness. This kernel captures the system’s spectral features, enabling SVMs to model complex physical phenomena with kernels reflecting underlying Green’s functions.
4.2.2 Constructing Custom Kernels via KPM
The Kernel Polynomial Method (KPM) provides a systematic approach to construct custom kernels that approximate Green’s functions. The general construction method is shown in Algorithm 2.
@@ 4.2.2 Constructing Custom Kernels via KPM
This construction makes sure that the resulting kernel is positive semi-definite, and therefore satisfies Mercer’s condition. For photonic crystals, additional periodicity might be incorporated by adjusting the polynomial expansion or combining it with periodic basis functions. This will further align the custom kernel with the system’s specific Green’s function.
@@
1:  Rescale $H$ to $\tilde{H}$ with eigenvalues in $[-1,1]$
@@
4.2.3 Example: Constructing a Photonic-Crystal Kernel via KPM
In Algorithm 2, we have introduced how the Kernel Polynomial Method (KPM) leverages orthogonal polynomials (e.g. Chebyshev polynomials) to construct kernels that remain positive semi-definite (PSD). We now illustrate how to apply that procedure (Algorithm 2) to produce a custom kernel for photonic crystals, thereby replacing any indefinite (sigmoid-like) kernel with a physically motivated alternative.
1) Identify the Physical Operator.
Photonic crystals are governed by the Helmholtz-type operator
@@ 4.2.3 Example: Constructing a Photonic-Crystal Kernel via KPM
2) Approximate via Orthogonal Polynomials.
Applying KPM (Algorithm 2) involves three main steps: (i) rescaling the operator $\hat{P}$ so that its spectrum lies within $[-1,1]$, (ii) expanding the resolvent or desired spectral function in a Chebyshev (or Jacobi) basis $\{T_{n}\}$, (iii) and damping the expansion coefficients $\mu_{n}$ to reduce Gibbs oscillations:
@@
Here, $\psi(\mathbf{x},\mathbf{x}^{\prime})$ is some function reflecting the distance or inner product between $\mathbf{x}$ and $\mathbf{x}^{\prime}$, typically $\mathbf{x}^{\top}\mathbf{x}^{\prime}$ after suitable normalization.
3) Embed Periodicity for Photonic Crystals.
To capture the photonic crystal’s periodic structure, we can combine the polynomial expansion with a periodic factor. For instance, set
$\psi(\mathbf{x},\mathbf{x}^{\prime})\;=\;\gamma\,\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}\quad\text{and}$
@@
$K_{\mathrm{crystal}}(\mathbf{x},\mathbf{x}^{\prime})\;=\;\Bigl{[}\sum_{n=0}^{N}\mu_{n}^{\prime}\;T_{n}\!\bigl{(}\psi(\mathbf{x},\mathbf{x}^{\prime})\bigr{)}\Bigr{]}\;\times\;\Pi\bigl{(}\mathbf{x},\mathbf{x}^{\prime}\bigr{)}.$
This kernel has two components: (a) a Chebyshev-based PSD expansion approximating the fundamental Green’s function decay, and (b) a cosine term reflecting the periodic modulation with period $p$.
In the photonic-crystal setting, we incorporate periodicity by multiplying the Chebyshev-based expansion with a term $\cos\!\bigl{(}\tfrac{2\pi}{p}\|\mathbf{x}-\mathbf{x}^{\prime}\|\bigr{)}$ to reflect the periodic refractive-index modulation. Practically, one must ensure that the overall kernel remains positive definite—e.g. by restricting the amplitude of the cosine term and verifying nonnegative expansions. This approach ensures that the kernel not only mirrors the decaying, wave-like nature of the photonic Green’s function but also remains PSD for stable SVM training.
4.2.4 Example: Constructing a PSD Kernel That Replaces Sigmoid-Like Kernels
We now give an explicit kernel formula, built from the Kernel Polynomial Method (KPM) (Algorithm 2), which can serve as a positive semi-definite (PSD) replacement for sigmoid-like kernels. Unlike the sigmoid kernel, whose positive semi-definiteness depends on restricted parameter ranges, this KPM-based kernel is guaranteed to remain PSD for all valid coefficient choices.
Step 1: Desired Functional Form.
Suppose we want a kernel $K_{\mathrm{desired}}(\mathbf{x},\mathbf{x}^{\prime})$ that roughly mimics the sigmoid’s sigmoidal response or saturating behavior, yet remains PSD. Instead of $\tanh(\gamma\,\mathbf{x}^{\top}\mathbf{x}^{\prime}+r)$, we choose to approximate a smooth “S-curve” by summing orthogonal polynomials with nonnegative expansion coefficients:
$K_{\mathrm{desired}}(\mathbf{x},\mathbf{x}^{\prime})\;\approx\;\sum_{n=0}^{N}\mu_{n}\,T_{n}\bigl{(}\varphi(\mathbf{x},\mathbf{x}^{\prime})\bigr{)},$
where $T_{n}$ are Chebyshev polynomials of the first kind, and $\varphi(\mathbf{x},\mathbf{x}^{\prime})$ is an inner product or distance-based function.
Step 2: Ensuring PSD via KPM.
We follow Algorithm 2:
  1. 1.
    Rescale domain: We define
    $\varphi(\mathbf{x},\mathbf{x}^{\prime})\;=\;\alpha\,\mathbf{x}^{\top}\mathbf{x}^{\prime}+\beta$
    so that the values lie in $[-1,1]$ (this typically requires a suitable normalization or bounding on $\mathbf{x}$).
  2. 2.
    Determine expansion coefficients: We pick a target function $f(z)$ that is smooth and sigmoidal on $z\in[-1,1]$, e.g. $f(z)\approx\tanh(\kappa z)$. Then we compute Chebyshev moments:
[…]
    and optionally apply a damping factor $g_{n}$ (Jackson kernel) to reduce Gibbs oscillations:
    $\mu_{n}^{\prime}\;=\;\mu_{n}\,g_{n}\,.$
  3. 3.
    Construct the kernel:
    $K_{\mathrm{KPM}}(\mathbf{x},\mathbf{x}^{\prime})\;=\;\sum_{n=0}^{N}\mu_{n}^{\prime}\;T_{n}\!\bigl{(}\alpha\,\mathbf{x}^{\top}\mathbf{x}^{\prime}+\beta\bigr{)}.$
    Because $\mu_{n}^{\prime}\geq 0$ and each $T_{n}$ term is integrated in a way consistent with Mercer’s theorem, $K_{\mathrm{KPM}}$ remains PSD.
Final Explicit Form.
Combining the above yields a fully explicit kernel:
$\boxed{K_{\mathrm{KPM}}(\mathbf{x},\mathbf{x}^{\prime})\;=\;\sum_{n=0}^{N}\Bigl{(}\mu_{n}\,g_{n}\Bigr{)}\,T_{n}\!\Bigl{(}\alpha\,\mathbf{x}^{\top}\mathbf{x}^{\prime}+\beta\Bigr{)}.}$ (32)
In practical SVM usage, one can tune $\alpha,\beta$ to ensure the argument stays within $[-1,1]$ for all data points, and choose the number of polynomial terms $N$ to balance accuracy with efficiency.
Applicability to Physical Problems.
Since $K_{\mathrm{KPM}}$ is PSD by construction, it can replace any indefinite/sigmoid kernel for:
  • Material classification or regression where a saturating (sigmoidal) response is desired, but $\tanh(\gamma\,\mathbf{x}^{\top}\mathbf{x}^{\prime}+r)$ fails Mercer’s condition.
  • Quantum many-body or PDE contexts needing kernels that softly saturate as $\|\mathbf{x}-\mathbf{x}^{\prime}\|$ grows (e.g. spin-lattice models). One can shape $f(z)$ to reflect the desired correlation decay.
  • Any domain (chemistry, biology, imaging) where indefinite kernels hamper SVM convergence. $K_{\mathrm{KPM}}$ supplies a stable, PSD alternative with tunable shape.
Because (32) is derived via orthogonal expansions on $[-1,1]$, it ensures Mercer-compatibility, thus sidestepping the instabilities of the sigmoid kernel.
Here, our main goal is to handle kernels such as $\tanh(\gamma\,\mathbf{x}^{\top}\mathbf{x}^{\prime}+r)$, which may become indefinite. By expanding a chosen saturating function $f(z)$ with nonnegative coefficients in a Chebyshev basis, we ensure the final kernel is PSD on $[-1,1]$. In practice, one rescales $\mathbf{x}^{\top}\mathbf{x}^{\prime}$ so the argument lies within $[-1,1]$. The resulting kernel $K_{\mathrm{KPM}}(\mathbf{x},\mathbf{x}^{\prime})$ inherits the smooth, saturating shape from $f(z)$ while satisfying Mercer’s condition, thus serving as a robust PSD substitute for the original sigmoid kernel.
4.3 Mismatch with Physical Green’s Functions
While some kernels (e.g., Gaussian/RBF) naturally align with Green’s functions under certain parameterizations, sigmoid-like kernels frequently fail to maintain the required positivity and Gaussian decay. In quantum mechanical and PDE contexts, Green’s functions are typically positive (or at least non-negative), ensuring well-defined probability amplitudes and correlation properties.
Green’s functions describe fundamental propagation and decay behaviors in physical systems. They frequently exhibit Gaussian or exponential-type decay in imaginary time (or distance), ensuring positivity in contexts like electron propagation. The sigmoid function, $\tanh(z)$, does not emulate this exponential decay. Instead, it saturates at $+1$ and $-1$, lacking the typical positivity constraints or the required “peak” shape of many Green’s functions.
Hence, in summary, by applying KPM, we circumvent the fundamental causes of the sigmoid kernel’s failure, namely, lack of guaranteed positivity, lack of guaranteed positive semidefiniteness, and a mismatch with typical Green’s function forms.
@@ 5 Conclusion and Future Outlook
In this paper, we have demonstrated a mathematical connection between Green’s functions and SVM kernels. We have furthermore shown that this connection leads to significant improvements in the predictive accuracy of SVM kernels arising from their correspondence with particular Green’s functions.
The kernels we tested were: the linear kernel, the RBF kernel, the polynomial kernel, and the sigmoid kernel. The sigmoid kernel proved to be an exception in the sense that it failed to capture the necessary characteristics of quantum mechanical phenomena. The reason for that was that the sigmoid kernel is positive semi-definite under certain conditions only. In cases where it is not positive semi-definite, it violates Mercer’s Theorem, under which only positive semi-definite kernels ensure convergence. The sigmoid kernel then inspired us to incorporate KPM into our custom kernel design, which facilitated the construction of a polynomial approximation of positive semi-definite functions. Such functions are necessary for convexity and, thereby, for good convergence. Hence, employing KPM for kernel design enables the development of custom kernels that approximate Green’s functions, enhancing the SVM’s ability to model complex systems.
It is our hope that this interdisciplinary framework will open new avenues for applying machine learning techniques to complex quantum mechanical problems.
@@ Acknowledgments
The authors wish to express their sincere gratitude to Professor Chong-Der Hu of National Taiwan University for his valuable suggestions and insightful discussions on many-body physics, which greatly enhanced the quality of this work. Renata Wong acknowledges support from the National Science and Technology Council grant No. NSTC 114-2112-M-182-002-MY3.
@@
The accompanying Python code includes the following:
  1. 1.
    Computing Conductivity using the RBF Kernel. (svm_conductivity.py)
  2. 2.
    Scattering Amplitudes using the Linear Kernel. (scattering_linear_kernel.py)
  3. 3.
    Anharmonic Oscillator Energy Levels using the Polynomial Kernel.
    (anharmonic_polynomial_kernel.py)
  4. 4.
    Phonon Dispersion using a Custom Periodic Kernel. (custom_kernel.py)
and can be accessed on the cloud server: https://drive.google.com/drive/folders/16BhLyXQogK1oLoOQMJyp5KkV6d9cSOlS.

### v2

@@ 5 Custom kernel design with KPM
5.1 Handling non-PSD kernels in SVMs
To address the limitations of non-PSD kernels, several strategies have been proposed, such as spectral correction [16, 23], pseudo-Euclidean embedding [21, 15], difference of convex functions (DC) programming [17], and iterative methods [25].
Spectral correction methods involve clipping, flipping, or shifting eigenvalues. These techniques can modify the kernel matrix to enforce positive semi-definiteness. Despite their simplicity, spectral correction methods can result in the loss of crucial structural information embedded in the original kernel. Pseudo-Euclidean embedding involves a set of methods for embedding the data into Krein spaces or pseudo-Euclidean spaces to handle indefinite kernels. Although effective, these methods increase computational complexity and are less interpretable. DC programming reformulates SVM optimization problems into convex and concave parts. This approach ensures convergence to a locally optimal solution but may involve higher computational overhead.
Here, we propose a custom kernel construction based on the KPM. Such a custom kernel can be used whenever there is no close correspondence between a kernel and a Green’s function for a given problem. KPM does not guarantee positive semidefiniteness. But it can be achieved to some degree by applying Jackson kernel smoothing to dampen Gibbs oscillations that may cause negative values, and post-processing. A post-processing step involves some sort of a correction. One straightforward correction method one could apply is for example the spectral correction mentioned at the beginning of this section, or a projection to the nearest PSD matrix. In this, one must ensure that the alignment with the associated Green’s function is only minimally distorted.
5.2 KPM-based kernel construction
@@
where $T_{n}(x)$ are Chebyshev polynomials, and $\mu_{n}$ are expansion coefficients. To ensure stability and convergence, the spectrum of $A$ is scaled to [1,1][-1,1][ - 1 , 1 ] and damping kernels (e.g., Jackson kernel) adjust the coefficients:
@@
The general construction method is shown in Algorithm 1.
@@
    If needed, post-process to make the kernel positive semidefinite.
  5. 5.
@@
1: Rescale $H$ to $\tilde{H}$ with eigenvalues in [1,1][-1,1][ - 1 , 1 ]
@@
5.3 Example: constructing a photonic crystal kernel via KPM
In Algorithm 1, we have introduced how KPM makes use of orthogonal polynomials to construct custom kernels that align with Green’s functions. We now illustrate how to apply that procedure to produce a custom kernel for photonic crystals.
Step 1: Identify the physical operator. Photonic crystals are governed by the Helmholtz-type operator
@@ 5.3 Example: constructing a photonic crystal kernel via KPM
Step 2: Approximate via orthogonal polynomials. Applying Algorithm 1 involves three main steps: (i) rescaling the operator $\hat{P}$ so that its spectrum lies within [1,1][-1,1][ - 1 , 1 ], (ii) expanding the resolvent or desired spectral function in a Chebyshev basis $\{T_{n}\}$, and (iii) damping the expansion coefficients $\mu_{n}$ to reduce Gibbs oscillations:
@@
Here, $\psi(\mathbf{x},\mathbf{x}^{\prime})$ is some function reflecting the distance or inner product between $\mathbf{x}$ and $\mathbf{x}^{\prime}$. For the Chebyshev expansion to be valid, the argument of the polynomials must be scaled to the interval [1,1][-1,1][ - 1 , 1 ]. Therefore, we define a rescaled function, for example, $\psi^{\prime}(x,x^{\prime})=a\psi(x,x^{\prime})+b$, where the constants $a$ and $b$ are chosen such that $\psi^{\prime}$ maps the relevant range of distances for the dataset into this interval. The expansion is then performed on this rescaled variable.
Step 3: Embed periodicity for photonic crystals. To capture the photonic crystal’s periodic structure, we can combine the polynomial expansion with a periodic factor. For instance, set
$\psi(\mathbf{x},\mathbf{x}^{\prime})\;=\;\gamma\,\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}\quad$
and
@@
As $\Pi(\mathbf{x},\mathbf{x}^{\prime})$ is not always positive due to the oscillatory nature of the cosine term, it needs to be replaced with a non-negative variant such as
$\Pi^{+}(\mathbf{x},\mathbf{x}^{\prime})=\frac{1}{2}\Big{(}1+\cos\Big{(}\tfrac{2\pi}{p}\,\|\mathbf{x}-\mathbf{x}^{\prime}\|\Big{)}\Big{)}$
which maps the cosine output from [1,1][-1,1][ - 1 , 1 ] to [0,1][0,1][ 0 , 1 ] while retaining its periodic structure.
@@
$K_{\mathrm{crys}}(\mathbf{x},\mathbf{x}^{\prime})\;=\;\Bigl{[}\sum_{n=0}^{N}\mu_{n}^{\prime}\;T_{n}\!\bigl{(}\psi(\mathbf{x},\mathbf{x}^{\prime})\bigr{)}\Bigr{]}\;\times\;\Pi^{+}\bigl{(}\mathbf{x},\mathbf{x}^{\prime}\bigr{)}.$
In practice, one must ensure that the overall kernel remains positive definite, e.g. by restricting the amplitude of the cosine term and verifying nonnegative expansions.
@@ 6 Conclusion and future outlook
In this paper, we have demonstrated a mathematical connection between Green’s functions (and, by extension, quantum propagators) and SVM kernels. We have furthermore shown that this connection leads to significant improvements in the predictive accuracy of SVM kernels arising from their correspondence with particular Green’s functions.
The kernels we tested were: the linear kernel, the RBF kernel, the polynomial kernel, and the sigmoid kernel. The sigmoid kernel proved to be an exception in the sense that it failed to capture the necessary characteristics of quantum mechanical phenomena. The reason for that was that the sigmoid kernel is positive semi-definite only under certain conditions. In cases where it is not positive semi-definite, it violates Mercer’s theorem, under which only positive semi-definite kernels ensure convergence. We then proposed a custom construction method for kernels to align with Green’s functions. The limitation of the method is that it doesn’t always guarantee that the resulting kernel is positive semidefinite. In the cases where it is not, one must additionally apply some correction techniques to approximate the kernel to a positive semidefinite matrix.
@@ Acknowledgments
The authors wish to express their sincere gratitude to Professor Chong-Der Hu of National Taiwan University and Assistant Professor Tsung-Wei Chiang of National Chung Cheng University for their valuable suggestions and insightful discussions on many-body physics, which greatly enhanced the quality of this work. Renata Wong acknowledges support from the National Science and Technology Council grant No. NSTC 114-2112-M-182-002-MY3 and from Chang Gung Memorial Hospital grant No. BMRPL94.
@@
Source code is available at GitHub repository https://github.com/kuonanhong/Code Available/Quantum Journal-Kernel and Green Function Correspondence. The code is organized into modules corresponding to each application:
  • svm_conductivity.py for electrical conductivity using the RBF kernel.
  • scattering_linear_kernel.py for electron-proton scattering using the linear kernel.
  • anharmonic_polynomial_kernel.py for anharmonic oscillator energy levels using polynomial kernels.
  • custom_kernel.py for photonic crystal modeling using a custom periodic kernel.
The repository includes detailed instructions for reproducibility and examples of parameter tuning.